OpenStudy (anonymous):

Is 36 a perfect square? Explain

OpenStudy (anonymous):

yes its root is 6

hero (hero):

Hint: \(6 \times 6 = 36 \)

OpenStudy (lgbasallote):

@higgs not exactly. it's \(\pm 6\)

OpenStudy (lgbasallote):

it's not just 6. 6 is just the principal square root

OpenStudy (anonymous):

im still not getting how 26 is a perfict sq cuz its root is 6..

OpenStudy (lgbasallote):

you wrote 36

OpenStudy (lgbasallote):

26 is not a perfect square

OpenStudy (anonymous):

ik i put 26 by misstake i ment to put 36

OpenStudy (anonymous):

-6 x -6 = 36 6 x 6 = 36

hero (hero):

2nd Hint: \(6^2 = 36 \)

OpenStudy (lgbasallote):

\[\sqrt{36} = \sqrt{6^2} = 6\]

OpenStudy (anonymous):

or -6

OpenStudy (anonymous):

ok thanks

OpenStudy (lgbasallote):

wait let me rephrase that \[\sqrt{36} = \sqrt{6^2} = (6^2)^{1/2} = 6^{2/2} = 6^1 = 6\]

OpenStudy (lgbasallote):

or -6 yeah

OpenStudy (anonymous):

Now you ask, why is 36 a perfect square. Well, because, you can arrange 36 unit squares(squares of side 1 each) to form a big square.